OpenStudy (truwhovian11):

MEDAL AND FAN FOR THE FIRST PERSON TO HELP ME cristofer's baseball team has won 3/10 of its games while rodgers team has won 1/6 of its games which team won a greater fraction of its games

OpenStudy (truwhovian11):

@NetflixAndNo

OpenStudy (truwhovian11):

@vsvampires309

OpenStudy (truwhovian11):

@Virnios

OpenStudy (virnios):

I think 3/10 is better than 1/6.

OpenStudy (virnios):

if im wrong ill kill myself

OpenStudy (texaschic101):

3/10 = 18/60 1/6 = 10/60 yep....3/10 is bigger

OpenStudy (truwhovian11):

THANKS @Virnios ur correct